OpenStudy (anonymous):

what is the maximum value of xy when x+y =3

OpenStudy (anonymous):

xy=x(3-x)=3x-x^2 differentiating this will give 3-2x in maximum differentiation is zero 3-2x=0 x=1/2

OpenStudy (anonymous):

and answer is ?

OpenStudy (anonymous):

you know x, substitute it to x+y=3 and find y and then multiply x and y

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

(y)

OpenStudy (phi):

that idea word, except for the last step 3-2x=0 you get 3= 2x and x= 3/2 (not ½)

OpenStudy (anonymous):

yes. sorry, my bad

OpenStudy (anonymous):

Mathematica confirms phi's answer. Refer to the attachment where Maximize is a function from Mathematica's Constrainted Optimization function suite.